OpenStudy (anonymous):

list the values for x after completing the square: x^2+2x-143=0

OpenStudy (anonymous):

x^2+2x+1-1-143=0 (x+1)^2-144=0 (x+1)^2=144 x+1=+-12 x=11, x=-13

OpenStudy (amistre64):

(x+1)^2 = 143 +1 x = -1 +- sqrt(144) x = -1 +- 12; x = -13, x=11